Circular No. 8182

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

S/2003 (379) 1 J. L. Margot, California Institute of Technology, reports the discovery

on Aug. 14.6 UT of a satellite of (379) Huenna (V = 12.9) from adaptive optics H-band imaging with the 10-m W. M. Keck II telescope on Mauna Kea. The satellite was observed to follow the primary at the non-sidereal rate of roughly 10′′/hr for a period exceeding 2 hr. On Aug. 14.5518 UT, the companion was at a separation of 0′′.89 (projected separation 1200 km) in p.a. 34^{o}. The primary-to-secondary H-band flux ratio is ∼ 165, giving a diameter ratio of ∼ 13.

SUPERNOVAE 2003he AND 2003hf Further to IAUC 8179, J. Graham, H. Pugh, and W. Li report the

LOTOSS discovery of two apparent supernovae on unfiltered KAIT images:

SN  2003 UT  α_{2000}  δ_{2000}  Mag.  Offset
2003he Aug. 11.4 23^{h}58^{m}57.^{s}05 − 2^{o}14′53′′.1 18.7 2′′.5 W, 6′′.1 N
2003hf Aug. 14.2 16 50 48.38 +45 24 01.8 18.4 1′′.7 W, 6′′.1 S

Additional approximate KAIT magnitudes: SN 2003he in MCG –01-1-10, Aug. 6.4 UT, [19.5; 12.4, 18.1. SN 2003hf in UGC 10586, Aug. 10.2, [19.5; 15.2, 17.7.

COMET C/2003 H1 (LINEAR)

Improved parabolic orbital elements from MPC 49277:

Epoch = 2004 Feb. 5.0 TT

T = 2004 Feb. 22.6097 TT  ω = 196.^{o}1344
e = 0.999381  Ω = 18.9997 }2000.0
q = 2.239691 AU  i = 138.6689

COMET C/2001 Q4 (NEAT) Total visual magnitude estimates: July 22.69 UT, 13.0 (M. Mattiazzo,

Wallaroo, South Australia, 0.28-m reflector); 27.71, 12.9 (Mattiazzo); Aug. 4.70, 12.6 (D. A. J. Seargent, Cowra, N.S.W., 0.25-m reflector); 6.71, 12.5 (Seargent).

2003 August 15  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2003 CBAT  Daniel W. E. Green